OpenStudy (anonymous):

Shanee bikes 5 miles to the park on 30 minutes and 3 miles to the library in 45 minutes. What was her average speed?

OpenStudy (jack1):

5 miles @ 30 mins = ??? miles per hour? @helpme375

OpenStudy (anonymous):

.166666666667

OpenStudy (jack1):

then 3 miles @45 mins = ??? miles per hour?

OpenStudy (jack1):

ummm... no, 5 miles in 30 mins = 10 miles per hour

OpenStudy (jack1):

(30 mins = 1/2 of an hour)

OpenStudy (jack1):

and 3 miles in 45 mins = 1 mile every 15 mins 1 mile every 1/4 of an hour so = 4 miles per hour

OpenStudy (jack1):

so total time = 1/2 hour + 3/4 hour = 5/4 hours (1 hr 15 mins) and speed = 10 miles per hour for 30 / 75 mins and = 4 miles per hour for 45/75 mins = \[(10 \times \frac{ 30 }{ 75 } )+(4 \times \frac{ 45 }{ 75 }) = average speed\]